# Workflow as a Service Broker in Cloud Environment: A Systematic Literature Review

### v1

@@ Abstract.
Cloud computing has emerged as a promising platform for running scientific workflows across various domains. Scientists can take advantage of different cloud service models, such as serverful or serverless, to execute workflows based on their specific requirements, along with diverse pricing models like on-demand, reserved, or spot instances to reduce execution costs. However, the challenge of selecting appropriate resources and pricing models, coupled with the orchestration and scheduling of workflow tasks, creates significant complexity for users. To mitigate this burden, Workflow as a Service (WaaS) brokers have been introduced to facilitate workflow execution. In recent years, numerous studies have been published, either directly or indirectly related to this research area, highlighting the need for a comprehensive and systematic review of WaaS brokers to identify key trends and challenges in this field. In this paper, we conduct a Systematic Literature Review (SLR) on WaaS brokers within cloud environments. The SLR employs a thorough 3-tier strategy (database search, backward snowballing, and forward snowballing) to answer five research questions. A total of 74 high-quality articles, published in 43 prestigious venues, are analyzed to derive a taxonomy based on the architecture of WaaS brokers. The articles are classified and surveyed according to this taxonomy, and future research directions for the design and implementation of WaaS brokers are explored. This study provides valuable insights for researchers and developers, helping them identify major trends and issues in the field of WaaS brokers.
@@
Workflow as a Service, cloud broker, multiple workflow scheduling, systematic literature review
ccs: Computer systems organization Cloud computingccs: Theory of computation Scheduling algorithmsccs: Computing methodologies Distributed computing methodologiesccs: Computer systems organization Architectures Distributed architectures Cloud computing
1. Introduction
@@
Cloud computing has emerged as one of the leading technologies, transforming industries and reshaping the world. It allows customers to access and rent the resources they need on demand, paying on a pay-as-you-go basis instead of purchasing resources outright (Josep et al., 2010). Both industry and academia have embraced this new computational paradigm to meet their resource requirements (Versluis and Iosup, 2021). In particular, new paradigms in Artificial Intelligence, such as deep learning and large language models, can leverage large pools of resources, such as GPUs, to address their computational demands. Furthermore, leading cloud providers, such as Amazon, are striving to introduce new services to cater to evolving user needs and attract new customers. As a result, alongside traditional services like Infrastructure as a Service (IaaS), which provides Virtual Machines (VMs), we now have services such as Container as a Service (CaaS) and Function as a Service (FaaS), or serverless computing (Shafiei et al., 2022), which offer higher-level services beyond raw VMs.
Nowadays, many applications running on cluster and cloud resources take the form of workflows (Versluis and Iosup, 2021). The concept of workflows, originally rooted in the manufacturing and business process domains, has evolved into the scientific community, where it is now referred to as scientific workflows (Hilman et al., 2020). A scientific workflow describes a process aimed at achieving a scientific objective, typically expressed in terms of tasks and their dependencies. Scientific workflows are often represented as a Directed Acyclic Graph (DAG), where each vertex corresponds to a task, and each edge denotes a data or control dependency between two tasks. Researchers from various scientific domains commonly define their applications as scientific workflows (Juve et al., 2013; Callaghan et al., 2024).
Cloud computing provides an attractive environment for executing scientific workflows. Scientists can leverage various cloud services, such as IaaS, CaaS, or FaaS, alongside different pricing models, including on-demand, spot, or reserved, based on their specific needs and requirements, to execute workflows efficiently and cost-effectively. However, selecting appropriate resources and pricing models, and more critically, orchestrating and scheduling workflow tasks on these resources, poses a significant challenge for users. This is where cloud brokers come into play. A broker plays a vital role in addressing these challenges by mediating between cloud providers and customers to identify the most suitable services while considering user preferences (Li et al., 2022b). Additionally, brokers can assist users in avoiding the vendor lock-in problem, which refers to the dependence on a specific cloud provider’s services, potentially leading to monopolistic scenarios.
@@
Workflow as a Service (WaaS) is a type of cloud broker that accepts users’ workflows along with their specified Quality of Service (QoS) requirements, such as deadline and budget, and executes these workflows by provisioning and allocating appropriate resources from cloud provider(s). While the broker seeks to optimize its own benefits, it also aims to provide sufficient incentives for users to delegate their workflows, primarily by reducing the cost of execution. To achieve this, the broker operates through two key components: the resource provisioner and the task scheduler. The resource provisioner is tasked with selecting suitable resources from various service models (IaaS, SaaS, or FaaS) and pricing models (on-demand, spot, and reserved). For example, to minimize costs, the provisioner might rent a large VM and use containerization to share it among tasks from multiple workflows. Additionally, while most users rely on costly on-demand resources, the provisioner can opt for cheaper alternatives, such as spot instances, by managing their unreliability, or reserved instances for workflows that run over extended periods. The scheduler is responsible for mapping the tasks of individual workflows onto the provisioned resources. By running multiple workflows concurrently, the scheduler can optimize resource utilization more effectively than scheduling each workflow independently. Consequently, the scheduling challenge in WaaS brokers is often referred to as the Multiple Workflow (or Multi-Workflow) Scheduling Problem (MWSP).
@@
A comprehensive and systematic review of research on WaaS brokers is essential for identifying major trends and issues in this field. Such a review can serve as a valuable guideline for researchers and enthusiasts, providing a deeper understanding of the progress made so far and the challenges that remain to be addressed. To ensure proper coverage of research studies in the domain of WaaS brokers, a systematic methodology is critical for identifying all relevant search spaces and studies (Petersen et al., 2015; Kitchenham et al., 2009). To the best of our knowledge, no prior Systematic Literature Review (SLR) has been conducted specifically on WaaS brokers. While a few studies address the broader domain of cloud brokerage (Li et al., 2022b; Khorasani et al., 2024), they do not specifically examine WaaS brokers. However, given that the scheduler is a key component of a WaaS broker, surveys on workflow scheduling in cloud environments are considered related works. Several studies have surveyed cloud workflow scheduling in general (Versluis and Iosup, 2021; Adhikari et al., 2019; Rodriguez and Buyya, 2017), but while these articles touch on the Multiple Workflow Scheduling Problem (MWSP), they do not comprehensively address it. Recently, Saeeidzade and Ashtiani (Saeedizade and Ashtiani, 2024) published a survey on workflow scheduling in cloud environments, which includes a subsection on scheduling in WaaS brokers. However, this survey does not comprehensively explore the domain. The most relevant work to this SLR is by Hilman et al. (Hilman et al., 2020), which provides a detailed survey of MWSP in multi-tenant distributed systems such as clusters, grids, and clouds. While this survey is comprehensive, it has several limitations. First, its taxonomy encompasses all distributed systems, including aspects irrelevant to the cloud environment. Second, it was published in 2019 and reviews only 31 articles, whereas numerous new studies have emerged in the past five years, particularly in the field of WaaS brokers (as shown in Section 3). This highlights the need for an updated survey. Finally, new technologies, such as serverless computing, have gained significant attention since the publication of Hilman et al.’s survey, further underscoring the necessity for a new review focused on recent developments in this area.
@@
We have conducted a systematic literature review (SLR) (Kitchenham et al., 2009) in the field of WaaS brokers within the cloud environment. To identify studies related to this area, we designed queries based on key terms relevant to the field. A three-tier search strategy was employed, consisting of database search, backward snowballing, and forward snowballing (see Section 3). By integrating these three search strategies, 74 papers from 43 different venues (journals and conferences) were selected based on specific inclusion and exclusion criteria. These papers were then analyzed to answer 5 research questions, serving as a valuable resource for research teams and developers in the field. All included papers and the extracted information for answering the research questions are available in an Excel supplementary file, accessible at https://github.com/SaeidAbrishami/WaaSBroker.
@@
The remainder of this article is organized as follows: Section 2 provides a detailed introduction to the architecture of WaaS brokers. Section 3 elaborates on the research methodology and provides a brief report of the review. Section 4 proposes a taxonomy for this field and surveys existing solutions based on this taxonomy. The future directions of WaaS brokers are outlined in Section 5, and Section 6 concludes the article.
2. The architecture of WaaS broker
Scientific workflows are a common method for automating and executing complex scientific applications across various fields (Juve et al., 2013; Callaghan et al., 2024). These sophisticated applications typically require execution on resource-rich computing environments, such as distributed infrastructures. Cloud computing offers an ideal infrastructure for running these workflows by providing various services, including IaaS, CaaS, and FaaS, with different pricing models like on-demand, spot, and reserved. However, selecting the appropriate services and resources, along with the correct pricing model, poses a significant challenge for scientists as users. Additionally, efficiently utilizing resources through optimal scheduling of workflow tasks is another challenge for users. Workflow as a Service is an emerging paradigm that allows users (scientists) to execute their scientific workflows without worrying about resource management and task scheduling. WaaS is a type of broker that receives scientific workflows along with their required QoS, then manages the provisioning of necessary resources from cloud providers and schedules workflow tasks on these resources.
Figure 1. The architecture of a Workflow as a Service broker.
@@ 2. The architecture of WaaS broker
Figure 1 illustrates the architecture of a typical WaaS broker. As shown, there are three main actors in the system: users, cloud provider(s), and the WaaS broker. The process begins when users send one or more workflows to the WaaS broker. A workflow can be represented using a DAG, where each node corresponds to a task, and each edge represents a data or control dependency between two tasks. This dependency indicates that a child task can only begin execution once its parent task(s) has completed and the child task has received the necessary output data. Additionally, users can often specify QoS parameters, such as deadlines and maximum budgets, for executing the workflow through a Service Level Agreement (SLA). In some cases, the broker may agree to a (monetary) penalty if the SLA is violated.
The cloud provider offers the necessary services and resources to execute users’ workflows. Currently, three well-known services are provided: IaaS, CaaS, and FaaS. The traditional IaaS service provides VMs to users, with various hardware and software specifications. CaaS, a newer service, offers lightweight virtualized environments known as containers. Containers have become popular for managing applications in the cloud (Pahl et al., 2019). By using a container, a virtual environment is created with separate CPU, memory, block I/O, and network space, allowing multiple tasks to run in parallel on a single VM. These services are typically provided in two models. For example, Amazon Elastic Container Service (ECS) allows users to fully manage containers on their VMs leased through AWS EC2. However, ECS also offers a service called AWS Fargate, which enables users to run containers without worrying about where or on which VM they will be executed; Amazon itself manages the deployment and operation of these containers. Both types of containers have been employed by researchers. Finally, the most recent service provided by cloud providers is serverless computing, typically represented by the FaaS service (along with Backend as a Service (BaaS)). FaaS allows developers to write stateless, event-driven serverless functions, enabling them to focus on the logic of serverless applications without worrying about the resources (VMs or containers) on which the application will run; the provider manages these resources (Shafiei et al., 2022). AWS Lambda is the most well-known model of this service. Shared storage services, such as Amazon S3 and Amazon Elastic File System (EFS), are also essential for storing intermediate data between tasks.
@@
Another important aspect of the services of the cloud provider is the pricing model. On-demand is the oldest and most well-known pricing model: the cloud provider offers various services at pre-determined prices, and the user can lease resources on a pay-as-you-go basis. This means the user launches an instance of a resource whenever needed and terminates it when the task is complete, paying only for the time the resource was used, without any long-term commitment. Historically, users were charged by the hour, but now, major providers such as Amazon charge users per second. In contrast, the reserved pricing model allows the user to lease resources for an extended period (typically one year or more) at a significant discount compared to on-demand resources (around 40% off in Amazon). Finally, the spot pricing model enables users to access the cloud provider’s excess computing capacity at a substantial discount (up to 80% in Amazon). However, these resources are highly unreliable, as the cloud provider can reclaim them with as little as two minutes’ notice. Therefore, the user must plan accordingly to manage this uncertainty (Lin et al., 2022). The final pricing model pertains to the FaaS model, in which the user is typically billed based on the number of function invocations, the time it takes for execution, and the computational resources allocated to the function instance, particularly memory.
@@
Figure 1 shows that the WaaS broker consists of four components: negotiation and pricing, task scheduler, resource provisioner, and monitoring and fault tolerance. The negotiation and pricing component interacts with the user. It receives the user’s workflow along with its required QoS and determines if it can meet the requirements. If the answer is yes, it specifies the price for running the workflow. It is reasonable for the price to be lower than what the user would pay by directly leasing resources from the provider. However, the WaaS broker must also generate a profit. This can be achieved by efficiently using leased resources, sharing them between different workflows, and utilizing pricing models such as reserved instances to reduce the cost of workflow execution.
The task scheduler is responsible for mapping and arranging the workflow’s tasks onto the leased resources, optimizing certain criteria (such as makespan or execution cost) while meeting the workflow’s QoS requirements. To make efficient use of the leased resources, the scheduler should consider all tasks from different workflows simultaneously, rather than scheduling each workflow’s tasks individually. Therefore, the scheduler falls under the domain of Multi-Workflow Scheduling Problem (MWSP). Since the scheduler is the core component of the broker, as discussed in Section 3.2, these terms have been used as search strings to identify related articles. However, we have focused on articles that address the MWSP on leased resources (VMs or containers) from the client-side (or broker-side) and excluded papers that explore this problem from the provider-side, such as those that involve scheduling tasks on physical machines, such as (Wang et al., 2023, 2024; Fan et al., 2023a).
The resource provisioner is responsible for leasing appropriate resources based on the needs of the task scheduler. This component determines when, how many, and which type (in terms of computational power, such as the number of cores and memory) of resources should be requested from the provider, and when they should be released to avoid wasting money. It is important to note that the provisioner is not relevant in FaaS platforms, as resource management is handled by the provider itself. The final component is the monitoring and fault tolerance module, which is responsible for overseeing the correct execution of tasks and the status of the leased resources. Additionally, in the event of a failure, this component must take necessary actions to recover from the issue, such as re-executing the failed tasks. This is especially crucial in the spot pricing model, where resources have low reliability and can be reclaimed by the provider at short notice.
3. Research methodology

### v2

# Workflow as a Service Broker in Cloud Environment: A Systematic Mapping Study

@@ Abstract
Cloud computing has emerged as a promising platform for running scientific workflows across various domains. Scientists can take advantage of different cloud service models, such as serverful or serverless, to execute workflows based on their specific requirements, along with diverse pricing models like on-demand, reserved, or spot instances to reduce execution costs. However, the challenge of selecting appropriate resources and pricing models, coupled with the orchestration and scheduling of workflow tasks, creates significant complexity for users. To mitigate this burden, Workflow as a Service (WaaS) brokers have been introduced to facilitate workflow execution. In recent years, numerous studies have been published, either directly or indirectly related to this research area, highlighting the need for a comprehensive and systematic review of WaaS brokers to identify key trends and challenges in this field. In this paper, we conduct a Systematic Mapping Study (SMS) on WaaS brokers within cloud environments. The SMS employs a thorough 3-tier strategy (database search, backward snowballing, and forward snowballing) to answer five research questions. A total of 87 high-quality articles, published in 49 prestigious venues, are analyzed to derive a taxonomy based on the architecture of WaaS brokers. The articles are classified and surveyed according to this taxonomy, and future research directions for the design and implementation of WaaS brokers are explored. This study provides valuable insights for researchers and developers, helping them identify major trends and issues in the field of WaaS brokers.
keywords:
@@
Workflow as a Service, cloud broker, multiple workflow scheduling, systematic mapping study
journal: Computer Science Review
[…]
{highlights}
Conducts the first systematic mapping study on Workflow-as-a-Service brokers.
Develops a structured taxonomy for WaaS broker components and architectures.
Analyzes 87 primary studies to reveal prevailing trends and open research issues.
Outlines future research toward AI-driven, serverless, and hybrid quantum/classical brokers.
1 Introduction
@@
Cloud computing has become one of the most transformative technologies of the modern era, reshaping industries and revolutionizing how computing resources are accessed and utilized. It enables customers to rent resources on demand and pay on a pay-as-you-go basis, eliminating the need for large upfront investments in infrastructure [85]. Both industry and academia have widely adopted this computational paradigm to meet their growing and diverse resource requirements [169]. In particular, emerging paradigms in artificial intelligence (AI)—such as deep learning and large language models—benefit from the cloud’s ability to provide scalable access to resource-intensive infrastructure like GPUs. In response to evolving user demands, leading cloud service providers (CSPs) continue to innovate by offering new services aimed at improving usability, flexibility, and performance. As a result, beyond traditional offerings like Infrastructure as a Service (IaaS), which provides virtual machines (VMs), a range of higher-level services has emerged. These include Container as a Service (CaaS), Function as a Service (FaaS)—also known as serverless computing [146]—and Quantum as a Service (QaaS), also referred to as Quantum Computing as a Service (QCaaS) [95], each designed to abstract away infrastructure management and streamline application deployment.
Scientific workflows are widely used to automate and execute complex scientific applications across diverse domains [86, 2]. A scientific workflow defines a process aimed at achieving a scientific objective, typically represented as a set of tasks and their dependencies. To support the composition, planning, orchestration, and automated execution of such workflows, researchers commonly rely on Workflow Management Systems (WMSs) [158]—such as Pegasus [46], Apache Airavata [56], and Apache Airflow [57]. These applications usually require resource-rich computing environments, such as distributed infrastructures. Cloud computing offers an attractive solution, enabling scientists to exploit services such as IaaS, CaaS, and FaaS, together with pricing models like on-demand, spot, or reserved, to execute workflows efficiently and cost-effectively. A central challenge, however, lies in selecting appropriate resources and pricing models and, more critically, orchestrating and scheduling workflow tasks across them. This is where cloud brokers become essential. Brokers mediate between CSPs and customers, identifying suitable services while incorporating user preferences [90]. They also mitigate the vendor lock-in problem, i.e., dependence on a single CSP’s services, which may otherwise lead to monopolistic scenarios.
@@
Workflow as a Service (WaaS) represents a class of cloud brokers that accept users’ workflows together with their specified Quality of Service (QoS) requirements, such as deadline and budget, and execute them by provisioning and allocating suitable resources from one or more CSPs. While the broker seeks to maximize its own benefits, it must also provide sufficient incentives for users to outsource their workflows, most notably by reducing execution costs. To accomplish this, the broker relies on two primary components: the resource provisioner, which is responsible for selecting resources across different service and pricing models, and the task scheduler, which maps workflow tasks onto the provisioned resources. By executing multiple workflows concurrently, the scheduler can improve overall resource utilization compared to scheduling workflows independently. This makes scheduling in WaaS brokers a distinctive challenge, commonly referred to as the Multiple Workflow (or Multi-Workflow) Scheduling Problem (MWSP).
A structured review of WaaS brokers is crucial for uncovering trends, challenges, and gaps in this domain. To achieve such coverage, systematic methodologies are typically adopted, most notably the Systematic Literature Review (SLR) [92] and the Systematic Mapping Study (SMS) [127]. While both approaches share a rigorous procedure for locating, filtering, and selecting studies, their objectives are fundamentally different. An SLR is a focused, evidence-oriented method designed to synthesize and compare existing research in depth, often including a critical evaluation of methods, metrics, and performance results under similar conditions. By contrast, an SMS does not aim to compare or rank methods. Instead, its role is broader and more exploratory: it maps the landscape of existing work, categorizes studies by dimensions such as method, contribution type, or evaluation strategy, and analyzes descriptive statistics such as publication years, venues, and frequency of techniques. The emphasis is on identifying research trends and highlighting underexplored areas, thereby revealing gaps in the literature. In this sense, SMS research questions are intentionally general, seeking to provide a big picture of the field, whereas SLRs drill down into fine-grained, evidence-based comparisons. An SMS often lays the groundwork for one or more SLRs by highlighting sub-areas that merit detailed comparative study.
@@
To the best of our knowledge, no prior SMS has been conducted specifically on WaaS brokers. While a few studies address the broader domain of cloud brokerage [106, 90], they do not specifically examine WaaS brokers. However, given that the scheduler is a key component of a WaaS broker, surveys on workflow scheduling in cloud environments are considered related works. Several studies have surveyed cloud workflow scheduling in general [169, 3, 94, 116], but while these articles touch on the MWSP, they do not comprehensively address it. Recently, Saeeidzade and Ashtiani [140] published a survey on workflow scheduling in cloud environments, which includes a subsection on scheduling in WaaS brokers. However, this survey does not comprehensively explore the domain. The most relevant work to this SMS is by Hilman et al. [68], which provides a detailed survey of MWSP in multi-tenant distributed systems such as clusters, grids, and clouds. While this survey is comprehensive, it has several limitations. First, its taxonomy encompasses all distributed systems, including aspects irrelevant to the cloud environment. Second, it was published in 2019 and reviews only 31 articles, whereas numerous new studies have emerged in the past five years, particularly in the field of WaaS brokers (as shown in Section 3). This highlights the need for an updated survey. Finally, new technologies, such as serverless and quantum computing, have gained significant attention since the publication of Hilman et al.’s survey, further underscoring the necessity for a new review focused on recent developments in this area.
@@
We have conducted an SMS [91, 127] in the field of WaaS brokers within the cloud environment. To identify studies related to this area, we designed queries based on key terms relevant to the field. A three-tier search strategy was employed, consisting of database search, backward snowballing, and forward snowballing (see Section 3). By integrating these three search strategies, 87 papers from 49 different venues (journals and conferences) were selected based on specific inclusion and exclusion criteria. These papers were then analyzed to answer 5 research questions, serving as a valuable resource for research teams and developers in the field. All included papers and the corresponding extracted data used to address the research questions are provided in Appendix B, as an Excel supplementary file available at https://github.com/SaeidAbrishami/WaaSBroker.
@@
The remainder of this article is organized as follows: Section 2 provides a detailed introduction to the architecture of WaaS brokers. Section 3 elaborates on the research methodology and provides a brief report of the review. Section 4 proposes a taxonomy for this field and surveys existing solutions based on this taxonomy. The future directions of WaaS brokers are outlined in Section 5, and Section 6 concludes the article. Furthermore, Appendix A discusses real-world adoption and industrial implementations of WaaS brokers.
2 Architecture of a WaaS broker
@@
Figure 1 illustrates the architecture of a typical WaaS broker, as proposed based on both academic literature [140, 3, 68, 163, 159, 190, 156, 165, 157, 110, 38, 125] and real-world implementations [46, 56]. As shown, there are three main actors in the system: users, cloud service provider(s), and the WaaS broker. The process begins when users send one or more workflows to the WaaS broker. A workflow is usually represented as a Directed Acyclic Graph (DAG), where nodes denote tasks and edges capture data or control dependencies between them. This dependency indicates that a child task can only begin execution once its parent task(s) has completed and the child task has received the necessary output data. Additionally, users can often specify QoS parameters, such as deadlines and maximum budgets, for executing the workflow through a Service Level Agreement (SLA). In some cases, the broker may agree to a (monetary) penalty if the SLA is violated.
CSPs—particularly the three major ones: Amazon Web Services (AWS), Google Cloud Platform (GCP), and Microsoft Azure—offer the essential services and resources required to execute users’ workflows. A broker can provision these resources from a single CSP, multiple CSPs, or a hybrid cloud environment that combines private (on-premise) and public cloud infrastructures (see Section 4.3.6). Currently, four widely used cloud services support the execution of scientific workflows: IaaS, CaaS, FaaS, and QaaS. Traditional IaaS provides users with VMs featuring various hardware and software configurations. However, the relatively long provisioning time of heavyweight VMs makes IaaS less suitable for small- to medium-sized tasks in traditional scheduling models, where each VM is typically assigned to a single task.
This limitation is addressed by CaaS, a newer model that offers lightweight virtualized environments known as containers. Containers have gained popularity for managing cloud applications due to their efficiency and flexibility [124]. Each container operates in an isolated environment with dedicated CPU, memory, block I/O, and network resources, enabling multiple tasks to run concurrently on a single VM. This allows task schedulers to place several small or medium-sized tasks on the same VM, achieving higher resource utilization compared to the IaaS model. Popular container services include Amazon Elastic Container Service (ECS), Azure Container Instances (ACI), and Google Kubernetes Engine (GKE).
Figure 1: Architecture of a Workflow as a Service broker.
Another recent service model offered by CSPs is serverless computing, typically realized through FaaS in combination with Backend-as-a-Service (BaaS). FaaS allows developers to write stateless, event-driven functions, enabling them to focus solely on application logic without managing the underlying infrastructure (e.g., VMs or containers), which is fully managed by the CSP [146]. Popular FaaS offerings include AWS Lambda, Azure Functions, and Google Cloud Functions. FaaS simplifies workflow execution by allowing tasks to be defined as individual functions, with resource provisioning and scaling handled automatically by the CSP. However, certain inherent characteristics of FaaS platforms introduce challenges. For instance, the stateless nature of functions requires the scheduler to rely on external storage services to store and exchange intermediate data between tasks. To address this, object storage solutions such as AWS S3, Azure Blob Storage, and Google Cloud Storage (GCS) are commonly used. When file-based access is required, shared file storage services—such as AWS Elastic File System (EFS), Azure Files, and Google Cloud Filestore—can be employed.
Recently, Quantum Computing (QC) has emerged as a revolutionary technology with the potential to transform numerous scientific domains by solving computational problems that are intractable for even the most advanced classical computers. Applications such as quantum chemistry, combinatorial optimization, and machine learning stand to benefit significantly from quantum speedups [58]. Due to the complexity and high cost of operating quantum hardware—along with strict environmental requirements—on-premise deployment is not practical for most users. As a result, quantum computing is typically offered through cloud-based platforms, a model known as QaaS [95]. Notable QaaS platforms include Amazon Braket [12], Azure Quantum [121], Google Quantum AI [66], IBM Quantum Platform [71], and IonQ [74]. These services give scientists and developers access to real quantum hardware and simulators, enabling experimentation and algorithm development without needing direct access to physical quantum computers.
Contemporary quantum applications are predominantly hybrid, involving both quantum and classical components. Typically, quantum tasks are executed on Quantum Processing Units (QPUs), while classical computing resources are used for pre-processing, post-processing, and coordination tasks [44]. A well-known class of such applications is Variational Quantum Algorithms (VQAs) [58, 45], which follow an iterative quantum–classical interaction model. In VQAs, a parameterized quantum circuit is evaluated by a quantum computer, and the resulting cost function is optimized using a classical algorithm such as gradient descent. This results in a feedback loop: the quantum device computes the current output, the classical optimizer updates the parameters based on the cost, and the updated parameters are passed back for re-execution—repeating until a stopping condition is met. These applications can be effectively modeled as hybrid scientific workflows, comprising a sequence of interconnected classical and quantum tasks [44, 45]. Executing such workflows requires seamless integration between High-Performance Computing (HPC) systems and QC environments [51, 24, 149]. The provisioning and scheduling of hybrid workflows introduce a new set of challenges, exceeding those found in purely classical systems. Addressing these challenges is essential for the development of robust and efficient WaaS brokers capable of supporting hybrid quantum-classical workloads [44].
@@
Another important aspect of the services of the CSPs is the pricing model. On-demand is the oldest and most well-known pricing model: the CSP offers various services at pre-determined prices, and the user can lease resources on a pay-as-you-go basis. Historically, users were charged by the hour, but now, major CSPs charge users per second. In contrast, the reserved pricing model allows the user to lease resources for an extended period (typically one year or more) at a significant discount compared to on-demand resources. Finally, the spot pricing model enables users to access the CSP’s excess computing capacity at a substantial discount. However, these resources are highly unreliable, as the CSP can reclaim them with as little as two minutes’ notice. Therefore, the user must plan accordingly to manage this uncertainty [108]. The final pricing model pertains to the FaaS model, in which the user is typically billed based on the number of function invocations, the time it takes for execution, and the computational resources allocated to the function instance, particularly memory.
@@
Figure 1 shows that the WaaS broker consists of four components: negotiation and pricing, task scheduler, resource provisioner, and monitoring and fault tolerance. The negotiation and pricing component interacts with the user. It receives the user’s workflow along with its required QoS and determines if it can meet the requirements. If the answer is yes, it specifies the price for running the workflow. It is reasonable for the price to be lower than what the user would pay by directly leasing resources from the CSP. However, the WaaS broker must also generate a profit. This can be achieved by efficiently using leased resources, sharing them between different workflows, and utilizing pricing models such as reserved instances to reduce the cost of workflow execution.
The task scheduler assigns workflow tasks to leased resources, aiming to optimize objectives such as makespan or execution cost while meeting QoS requirements. To ensure efficient resource usage, it considers tasks from all workflows jointly rather than scheduling each workflow in isolation. As a result, the scheduler directly addresses the MWSP. The resource provisioner complements the scheduler by leasing resources as needed. It decides when, how many, and which types of resources (e.g., cores, memory) to acquire, as well as when to release them to avoid unnecessary cost. This component is irrelevant in FaaS platforms, where resource management is handled directly by the CSP.
The fourth component, the monitoring and fault tolerance, ensures correct task execution and tracks the status of leased resources. In case of failures, it initiates recovery actions such as task re-execution—especially critical in spot pricing models, where resources are unreliable and may be reclaimed at short notice. Finally, the broker’s local storage supports workflow execution by maintaining input files (or container images) and metadata. This includes provenance data for reproducibility and monitoring data such as execution logs and resource usage records [158].
3 Research Methodology